# One Perturbation is Enough: On Generating Universal Adversarial Perturbations against Vision-Language Pre-training Models

### v3

@@ 1 Introduction
Adversarial attacks [7], which aim to deceive models during inference time, have attracted extensive attention due to their significant threat to security-critical scenarios [15]. Recent studies have shown that VLP models are also vulnerable to adversarial samples. The pioneering work Co-Attack [48] proposes the first multimodal attack that simultaneously perturbs both image and text modalities and displays excellent performance. However, Co-Attack only considers relatively easier white-box attacks where victim models are completely accessible. To handle more practical black-box settings, subsequent studies propose various transferable adversarial samples generated on an available surrogate model to fool other inaccessible models. Specifically, SGA [29] significantly improves the adversarial transferability through the set-level cross-modal guidance obtained from data augmentations. Subsequently, TMM [41] proposes to jointly destroy the modality-consistency features within the clean image-text pairs and include more modality-discrepancy features in the perturbations to further enhance transferability. While existing methods have achieved great success, they are all instance-specific and need to generate a perturbation for each input pair, which results in substantial computational overhead. Meanwhile, universal adversarial attacks (illustrated in Fig. 1), as an efficient instance-agnostic approach that uses only one Universal Adversarial Perturbation (UAP) to conduct attacks, have not been fully investigated for VLP models. This naturally leads to a question, is it possible to design a UAP that can deceive VLP models across different image-text pairs?
@@
Motivation. To this end, we make an intuitive attempt to transplant renowned approaches UAP [31] and GAP [37] to attack several VLP models by maximizing the distance between the embeddings of the adversarial image and its matched texts. Unfortunately, Fig. 2 demonstrates that these methods yield unsatisfactory attack success rates (ASR), especially for black-box attacks. Empirically, this failure stems from their narrow focus on the image modal, disregarding the other modality and the multimodal information that plays a pivotal role in VLP models. To overcome this challenge, we revisit VLP models’ basic training paradigm and emphasize that regardless of the downstream V+L tasks, their achieved outstanding performance is heavily reliant on the well-established multimodal alignment, which draws the embeddings of matched image-text pairs closer while distancing those of non-matched pairs. In light of this consideration, we argue that the core of an effective universal adversarial attack is to obtain a UAP that can fundamentally destroy this learned alignment relationship to mislead VLP models into making incorrect decisions. Besides, Fig. 2 shows that the generator-based GAP consistently outperforms UAP methods, due to the excellent distributional modeling capability of generators. This suggests the superiority of the generative paradigm, which is also corroborated by numerous studies [17, 16].
@@
Based on these insights, we propose a novel generative framework that learns a Contrastive-training Perturbation Generator with Cross-modal conditions (C-PGC) to launch universal attacks on VLP models. To essentially destroy the multimodal alignment, we devise to utilize VLP models’ most powerful weapons to attack against themselves. I.e., utilize the contrastive learning mechanism to train the generator based on maliciously constructed image-text pairs that completely violate the correct V+L matching relationship, to produce perturbation that pushes the embeddings of matched pairs apart while pulling those of non-matched ones together. Moreover, most previous studies [48, 29, 49] simply exploit the crucial cross-modal information by maximizing the feature distance between samples of different modals to optimize perturbations, without deeper exploration for attack enhancement. In contrast, we fully harness V+L characteristics by refining the generator architecture to incorporate cross-modal knowledge through cross-attention mechanisms for better guidance. Besides, we also consider the intra-modal influence and introduce a unimodal distance loss to further improve attack effectiveness. We highlight that the proposed framework is seamlessly compatible with text perturbation generation, achieving a truly multimodal universal attack that benefits from the synergy between V+L modalities. Our contributions are as follows:
@@ 2.1 Vision-Language Pre-training Models
VLP models are pre-trained on massive image-text pairs to learn the semantic correlations across modalities and serve diverse multimodal user demands [8, 13]. We next illustrate the basis of VLP models from multiple perspectives.
Architectures. Based on the ways of multimodal fusion, the architectures of VLP models can be classified into two types: single-stream and dual-stream architectures. Single-stream architectures [25, 9] directly concatenate the text and image features, and calculate the attention in the same Transformer block for multimodal fusion. On the contrary, dual-stream architectures [38, 24] separately feed the text and image features to different Transformer blocks and leverage the cross-attention mechanism for multimodal fusion. Generally, single-stream architectures are more parameter-efficient than dual-stream architectures since they adopt the same set of parameters in a Transformer block for the text and image modalities.
Pre-training Objectives. The pre-training objectives for VLP models mainly include masked feature completion, multimodal feature matching, and specific downstream objectives. Masked feature completion [9] encourages VLP models to predict the deliberately masked tokens based on the remaining unmasked tokens. Multimodal feature matching [23] pre-trains VLP models to precisely predict whether the given image-text pairs are matched. Specific downstream objectives [2] directly utilize the training objectives of downstream tasks (e.g., visual question answering) for pre-training VLP models.
Downstream Tasks. In this paper, we mainly consider the following multimodal downstream tasks: (1) Image-text retrieval (ITR) [42]: finding the most matched image for the given text and vice versa, including image-to-text retrieval (TR) and text-to-image retrieval (IR). (2) Image caption (IC) [4]: generating the most suitable descriptions for the given image. (3) Visual grounding (VG) [20]: locating specific regions in the image that correspond with the given textual descriptions. (4) Visual entailment (VE) [43]: analyzing the input image and text and predicting whether their relationship is entailment, neutral, or contradiction.
@@ 2.2 Adversarial Attacks
Instance-specific Attacks on VLP Models. The adversarial robustness of VLP Models has already become a research focus. Early works [22, 45] impose perturbations only on single modality and lack cross-modal interactions when attacking multimodal models. To address this issue, Co-Attack [48] conducts the first multimodal white-box attacks on VLP models. On the basis of Co-Attack, [29] extends the attacks to more rigorous black-box settings and proposes SGA, which utilizes set-level alignment-preserving argumentations with carefully designed cross-modal guidance. However, [41] points out that SGA fails to fully exploit modality correlation, and proposes TMM to better leverage cross-modal interactions by tailoring both the modality-consistency and modality-discrepancy features. Nonetheless, these methods are all instance-specific and need to craft perturbations for each input pair.
Universal Adversarial Examples. Universal adversarial attacks [31, 32, 51] aim to deceive the victim model by exerting a uniform adversarial perturbation to all samples. These attacks save the redundant procedures of redesigning perturbations for each input sample and are hence more efficient than instance-specific methods. Generally, universal adversarial attacks can be categorized into optimization-based [31, 40, 28] and generation-based [18, 17, 3, 51] methods. Benefiting from the powerful modeling abilities of generative models, generation-based methods are more versatile and can produce more natural samples than optimization-based ones. A concurrent work ETU [49] also investigates UAP on VLP models and proposes a data augmentation named ScMix. However, ETU adopts a non-generative approach that narrowly focuses on image UAP, failing to constitute a truly multimodal attack for VLP models. Moreover, ETU demonstrates insufficient attack effects, especially for black-box transferability. In contrast, we propose a generative multimodal attack framework based on a malicious variant of contrastive learning, which yields UAP with strong attack effects and high transferability.
@@ 3.1 Problem Statement
where denotes the operation that uses the output V+L features to obtain the final predictions, indicates the text perturbation strategy [48, 29] that replaces certain important tokens of the original sentence with crafted adversarial words, and is the correct prediction of the considered V+L task. To ensure the perturbation’s imperceptibility, we constrain the pixel-level image perturbation with norm of a given budget . Following previous studies [48, 29, 41], the textual perturbation is token-level and the stealthiness is accordingly constrained by the number of modified words . To ensure the stealthiness of text perturbation, we apply a rigorous restriction that permits only a single word to be substituted (). On the premise of imperceptibility, the attacker attempts to generalize the crafted UAP to a wider range of test data and victim models.
@@ 3.3 Detailed Design of C-PGC
Perturbation Generator Conditioned on Cross-modal knowledge. Previous generative universal attacks [17, 3] have shown excellent efficacy in fooling the discriminative models. Nevertheless, since existing generative attacks are limited to a single modality, directly utilizing the off-the-shelf generators might fail to leverage the multimodal interactions in these special V+L scenarios. To address this limitation, we additionally introduce cross-modal embeddings as auxiliary information to further facilitate the process of perturbation generation. Specifically, we modify the architecture of existing decoder-based generators by adding several cross-attention modules that have been proven effective in tasks with multiple input modalities. The obtained textual embeddings encoded by are then incorporated into our generator through:
@@
Multimodal Contrastive Loss. The preceding analysis regarding the failures of existing UAP attacks encourages us to design a loss function that can guide the generated UAP to break the learned multimodal feature alignment. Motivated by the fact that contrastive learning underpins the cross-modal alignment, we advocate leveraging this mechanism to attack VLP models themselves by contrastively training our C-PGC to essentially disrupt the benign alignment relationship. Concretely, we adopt the widely recognized InfoNCE [19] as our basic contrastive loss.
To establish the contrastive paradigm, we first define the adversarial image as the anchor sample. Besides, it is also necessary to construct an appropriate set of positive and negative samples. Based on the fundamental objective of our attack, it is natural to leverage the originally matched text description set as negative samples to amplify the discrepancy between matched image-text pairs in the feature space of VLP models. Additionally, we need to select a set of positive samples to further pull the adversarial image away from its corresponding text descriptions . To this end, we propose a farthest selection strategy, which associates the anchor image with target texts whose embeddings differ significantly from that of the original clean image , to reach a more strong disruption of the multimodal alignment. Specifically, we randomly sample a batch of text sets from and select the text set with the largest feature distances from the current image as positive samples, i.e., . Moreover, we utilize data augmentations that resize the clean into diverse scales and apply random Gaussian noise to acquire a more diverse image set for set-level guidance [29]. With these well-constructed positive and negative samples, the multimodal contrastive loss can be formulated as:
@@
A notable distinction between image and text attacks is the way to inject adversarial perturbations. We align with previous studies [48, 29, 41] and apply the token-wise substitute strategy that replaces certain important words in the original sentence with crafted adversarial words. Prior to the word replacement, a meticulous process is undertaken to identify the most optimal position within the sentence to insert the perturbation. Our strategy intends to replace the words that are more likely to have a greater influence during decision-making. Concretely, for each word within a given sentence, we compute the distance between the embeddings of the original sentence and the -masked version to determine its contribution. By convention, the imperceptibility of text UAP is controlled by the number of modified words [6]. As aforementioned, we set for high stealthiness, i.e., choose the single word exerting the highest feature distance as the target for replacement.
@@ 4.1 Experimental Setup
Downstream tasks and datasets. We conduct a comprehensive study of C-PGC on four downstream V+L tasks, including image-text retrieval (ITR), image captioning (IC), visual grounding (VG), and visual entailment (VE). For ITR tasks, we employ the Flickr30K [36] and MSCOCO [27] datasets which are commonly used in previous works [48, 29]. The MSCOCO is also adopted for evaluating the IC task. For VG and VE tasks, we evaluate on SNLI-VE [43] and RefCOCO+ [46] datasets respectively.
Models. We conduct experiments on a wide range of VLP models including ALBEF [23], TCL [44], X-VLM [47], CLIP [38], CLIP [38], and BLIP [24]. Note that for different V+L tasks, we correspondingly select different VLP models for evaluation based on their capability [41]. For instance, among the six considered VLP models, only ALBEF, TCL, and X-VLM can handle VG tasks, while only ALBEF and TCL can deal with VE tasks.
Baselines. We transplant the representative GAP [37] to V+L scenarios by appropriately editing its original loss function [29]. We also consider a concurrent UAP study ETU [49], which adopts a non-generative method that narrowly focuses on image perturbation, despite the multimodal nature of V+L scenarios. Note that it implements several versions and we report their best results.
Implementation details. Following the SGA [29], we adopt Karpathy split [21] to preprocess the dataset and build the test set for evaluation. The test set is disjoint with the generator’s training data for rigorous assessment. To ensure perturbation invisibility, we follow [41] and limit the perturbation budgets to and to . For augmentation, we resize the original images into five scales , and apply Gaussian noise . See Appendix B for more details.
@@ 4.2 Universal Attack Effectiveness
To align with previous studies [48, 29], we first consider the typical V+L task image-text retrieval and calculate the ASR as the proportion of successful adversarial samples within the originally correctly predicted pairs. We present the performance based on R@1 retrieval results in Table 1. Appendix C supplements the results of R@5 and R@10.
@@
White-box attack performance. By observing the white-box ASR in the gray-shaded area, we demonstrate that the proposed algorithm stably achieves excellent ASR on all the evaluated VLP models, validating the outstanding capability of the produced UAP. With only a single pair of perturbations, we reach a noteworthy average white-box ASR of nearly 90% on two large datasets in terms of both TR and IR tasks. Compared with GAP [37] and ETU [49], the proposed method consistently enhances the fooling rates in the white-box scenario, confirming the great validity of our suggested multimodal contrastive-learning mechanism. Essentially, the exceptional performance stems from the efficacy of our generated UAP in destroying the alignment between the image and text modalities, thereby misleading the VLP model during inference.

### v4

@@ 1 Introduction
Adversarial attacks [carlini2017towards], which aim to deceive models during inference time, have attracted extensive attention due to their significant threat to security-critical scenarios [eykholt2018robust]. Recent studies have shown that VLP models are also vulnerable to adversarial samples. The pioneering work Co-Attack [zhang2022towards] proposes the first multimodal attack that simultaneously perturbs both image and text modalities and displays excellent performance. However, Co-Attack only considers relatively easier white-box attacks where victim models are completely accessible. To handle more practical black-box settings, subsequent studies propose various transferable adversarial samples generated on an available surrogate model to fool other inaccessible models. Specifically, SGA [lu2023set] significantly improves the adversarial transferability through the set-level cross-modal guidance obtained from data augmentations. Subsequently, TMM [wang2024transferable] proposes to jointly destroy the modality-consistency features within the clean image-text pairs and include more modality-discrepancy features in the perturbations to further enhance transferability. While existing methods have achieved great success, they are all instance-specific and need to generate a perturbation for each input pair, which results in substantial computational overhead. Meanwhile, universal adversarial attacks (illustrated in Fig. 1), as an efficient instance-agnostic approach that uses only one Universal Adversarial Perturbation (UAP) to conduct attacks, have not been fully investigated for VLP models. This naturally leads to a question, is it possible to design a UAP that can deceive VLP models across different image-text pairs?
@@
Motivation. To this end, we make an intuitive attempt to transplant renowned approaches UAP [moosavi2017universal] and GAP [poursaeed2018generative] to attack several VLP models by maximizing the distance between the embeddings of the adversarial image and its matched texts. Unfortunately, Fig. 2 demonstrates that these methods yield unsatisfactory attack success rates (ASR), especially for black-box attacks. Empirically, this failure stems from their narrow focus on the image modal, disregarding the other modality and the multimodal information that plays a pivotal role in VLP models. To overcome this challenge, we revisit VLP models’ basic training paradigm and emphasize that regardless of the downstream V+L tasks, their achieved outstanding performance is heavily reliant on the well-established multimodal alignment, which draws the embeddings of matched image-text pairs closer while distancing those of non-matched pairs. In light of this consideration, we argue that the core of an effective universal adversarial attack is to obtain a UAP that can fundamentally destroy this learned alignment relationship to mislead VLP models into making incorrect decisions. Besides, Fig. 2 shows that the generator-based GAP consistently outperforms UAP methods, due to the excellent distributional modeling capability of generators. This suggests the superiority of the generative paradigm, which is also corroborated by numerous studies [gao2024nuat, feng2023dynamic].
@@
Based on these insights, we propose a novel generative framework that learns a Contrastive-training Perturbation Generator with Cross-modal conditions (C-PGC) to launch universal attacks on VLP models. To essentially destroy the multimodal alignment, we devise to utilize VLP models’ most powerful weapons to attack against themselves. I.e., utilize the contrastive learning mechanism to train the generator based on maliciously constructed image-text pairs that completely violate the correct V+L matching relationship, to produce perturbation that pushes the embeddings of matched pairs apart while pulling those of non-matched ones together. Moreover, most previous studies [zhang2022towards, lu2023set, zhang2024universal] simply exploit the crucial cross-modal information by maximizing the feature distance between samples of different modals to optimize perturbations, without deeper exploration for attack enhancement. In contrast, we fully harness V+L characteristics by refining the generator architecture to incorporate cross-modal knowledge through cross-attention mechanisms for better guidance. Besides, we also consider the intra-modal influence and introduce a unimodal distance loss to further improve attack effectiveness. We highlight that the proposed framework is seamlessly compatible with text perturbation generation, achieving a truly multimodal universal attack that benefits from the synergy between V+L modalities. Our contributions are as follows:
@@ 2.1 Vision-Language Pre-training Models
VLP models are pre-trained on massive image-text pairs to learn the semantic correlations across modalities and serve diverse multimodal user demands [chen2023vlp, du2022survey]. We next illustrate the basis of VLP models from multiple perspectives.
Architectures. Based on the ways of multimodal fusion, the architectures of VLP models can be classified into two types: single-stream and dual-stream architectures. Single-stream architectures [li2019visualbert, chen2020uniter] directly concatenate the text and image features, and calculate the attention in the same Transformer block for multimodal fusion. On the contrary, dual-stream architectures [radford2021learning, li2022blip] separately feed the text and image features to different Transformer blocks and leverage the cross-attention mechanism for multimodal fusion. Generally, single-stream architectures are more parameter-efficient than dual-stream architectures since they adopt the same set of parameters in a Transformer block for the text and image modalities.
Pre-training Objectives. The pre-training objectives for VLP models mainly include masked feature completion, multimodal feature matching, and specific downstream objectives. Masked feature completion [chen2020uniter] encourages VLP models to predict the deliberately masked tokens based on the remaining unmasked tokens. Multimodal feature matching [li2021align] pre-trains VLP models to precisely predict whether the given image-text pairs are matched. Specific downstream objectives [anderson2018bottom] directly utilize the training objectives of downstream tasks (e.g., visual question answering) for pre-training VLP models.
Downstream Tasks. In this paper, we mainly consider the following multimodal downstream tasks: (1) Image-text retrieval (ITR) [wang2016comprehensive]: finding the most matched image for the given text and vice versa, including image-to-text retrieval (TR) and text-to-image retrieval (IR). (2) Image caption (IC) [bai2018survey]: generating the most suitable descriptions for the given image. (3) Visual grounding (VG) [hong2019learning]: locating specific regions in the image that correspond with the given textual descriptions. (4) Visual entailment (VE) [xie2019visual]: analyzing the input image and text and predicting whether their relationship is entailment, neutral, or contradiction.
@@ 2.2 Adversarial Attacks
Instance-specific Attacks on VLP Models. The adversarial robustness of VLP Models has already become a research focus. Early works [kim2019single, yang2021defending] impose perturbations only on single modality and lack cross-modal interactions when attacking multimodal models. To address this issue, Co-Attack [zhang2022towards] conducts the first multimodal white-box attacks on VLP models. On the basis of Co-Attack, [lu2023set] extends the attacks to more rigorous black-box settings and proposes SGA, which utilizes set-level alignment-preserving argumentations with carefully designed cross-modal guidance. However, [wang2024transferable] points out that SGA fails to fully exploit modality correlation, and proposes TMM to better leverage cross-modal interactions by tailoring both the modality-consistency and modality-discrepancy features. Nonetheless, these methods are all instance-specific and need to craft perturbations for each input pair.
Universal Adversarial Examples. Universal adversarial attacks [moosavi2017universal, mopuri2018generalizable, zhou2023downstream] aim to deceive the victim model by exerting a uniform adversarial perturbation to all samples. These attacks save the redundant procedures of redesigning perturbations for each input sample and are hence more efficient than instance-specific methods. Generally, universal adversarial attacks can be categorized into optimization-based [moosavi2017universal, wang2023improving, liu2023enhancing] and generation-based [hayes2018learning, gao2024nuat, anil2024generating, zhou2023downstream] methods. Benefiting from the powerful modeling abilities of generative models, generation-based methods are more versatile and can produce more natural samples than optimization-based ones. A concurrent work ETU [zhang2024universal] also investigates UAP on VLP models and proposes a data augmentation named ScMix. However, ETU adopts a non-generative approach that narrowly focuses on image UAP, failing to constitute a truly multimodal attack for VLP models. Moreover, ETU demonstrates insufficient attack effects, especially for black-box transferability. In contrast, we propose a generative multimodal attack framework based on a malicious variant of contrastive learning, which yields UAP with strong attack effects and high transferability.
@@ 3.1 Problem Statement
where denotes the operation that uses the output V+L features to obtain the final predictions, indicates the text perturbation strategy [zhang2022towards, lu2023set] that replaces certain important tokens of the original sentence with crafted adversarial words, and is the correct prediction of the considered V+L task. To ensure the perturbation’s imperceptibility, we constrain the pixel-level image perturbation with norm of a given budget . Following previous studies [zhang2022towards, lu2023set, wang2024transferable], the textual perturbation is token-level and the stealthiness is accordingly constrained by the number of modified words . To ensure the stealthiness of text perturbation, we apply a rigorous restriction that permits only a single word to be substituted (). On the premise of imperceptibility, the attacker attempts to generalize the crafted UAP to a wider range of test data and victim models.
@@ 3.3 Detailed Design of C-PGC
Perturbation Generator Conditioned on Cross-modal knowledge. Previous generative universal attacks [gao2024nuat, anil2024generating] have shown excellent efficacy in fooling the discriminative models. Nevertheless, since existing generative attacks are limited to a single modality, directly utilizing the off-the-shelf generators might fail to leverage the multimodal interactions in these special V+L scenarios. To address this limitation, we additionally introduce cross-modal embeddings as auxiliary information to further facilitate the process of perturbation generation. Specifically, we modify the architecture of existing decoder-based generators by adding several cross-attention modules that have been proven effective in tasks with multiple input modalities. The obtained textual embeddings encoded by are then incorporated into our generator through:
@@
Multimodal Contrastive Loss. The preceding analysis regarding the failures of existing UAP attacks encourages us to design a loss function that can guide the generated UAP to break the learned multimodal feature alignment. Motivated by the fact that contrastive learning underpins the cross-modal alignment, we advocate leveraging this mechanism to attack VLP models themselves by contrastively training our C-PGC to essentially disrupt the benign alignment relationship. Concretely, we adopt the widely recognized InfoNCE [he2020momentum] as our basic contrastive loss.
To establish the contrastive paradigm, we first define the adversarial image as the anchor sample. Besides, it is also necessary to construct an appropriate set of positive and negative samples. Based on the fundamental objective of our attack, it is natural to leverage the originally matched text description set as negative samples to amplify the discrepancy between matched image-text pairs in the feature space of VLP models. Additionally, we need to select a set of positive samples to further pull the adversarial image away from its corresponding text descriptions . To this end, we propose a farthest selection strategy, which associates the anchor image with target texts whose embeddings differ significantly from that of the original clean image , to reach a more strong disruption of the multimodal alignment. Specifically, we randomly sample a batch of text sets from and select the text set with the largest feature distances from the current image as positive samples, i.e., . Moreover, we utilize data augmentations that resize the clean into diverse scales and apply random Gaussian noise to acquire a more diverse image set for set-level guidance [lu2023set]. With these well-constructed positive and negative samples, the multimodal contrastive loss can be formulated as:
@@
A notable distinction between image and text attacks is the way to inject adversarial perturbations. We align with previous studies [zhang2022towards, lu2023set, wang2024transferable] and apply the token-wise substitute strategy that replaces certain important words in the original sentence with crafted adversarial words. Prior to the word replacement, a meticulous process is undertaken to identify the most optimal position within the sentence to insert the perturbation. Our strategy intends to replace the words that are more likely to have a greater influence during decision-making. Concretely, for each word within a given sentence, we compute the distance between the embeddings of the original sentence and the -masked version to determine its contribution. By convention, the imperceptibility of text UAP is controlled by the number of modified words [behjati2019universal]. As aforementioned, we set for high stealthiness, i.e., choose the single word exerting the highest feature distance as the target for replacement.
@@ 4.1 Experimental Setup
Downstream tasks and datasets. We conduct a comprehensive study of C-PGC on four downstream V+L tasks, including image-text retrieval (ITR), image captioning (IC), visual grounding (VG), and visual entailment (VE). For ITR tasks, we employ the Flickr30K [plummer2015flickr30k] and MSCOCO [lin2014microsoft] datasets which are commonly used in previous works [zhang2022towards, lu2023set]. The MSCOCO is also adopted for evaluating the IC task. For VG and VE tasks, we evaluate on SNLI-VE [xie2019visual] and RefCOCO+ [yu2016modeling] datasets respectively.
Models. We conduct experiments on a wide range of VLP models including ALBEF [li2021align], TCL [yang2022vision], X-VLM [zeng2022multi], CLIP [radford2021learning], CLIP [radford2021learning], and BLIP [li2022blip]. Note that for different V+L tasks, we correspondingly select different VLP models for evaluation based on their capability [wang2024transferable]. For instance, among the six considered VLP models, only ALBEF, TCL, and X-VLM can handle VG tasks, while only ALBEF and TCL can deal with VE tasks.
Baselines. We transplant the representative GAP [poursaeed2018generative] to V+L scenarios by appropriately editing its original loss function [lu2023set]. We also consider a concurrent UAP study ETU [zhang2024universal], which adopts a non-generative method that narrowly focuses on image perturbation, despite the multimodal nature of V+L scenarios. Note that it implements several versions and we report their best results.
Implementation details. Following the SGA [lu2023set], we adopt Karpathy split [karpathy2015deep] to preprocess the dataset and build the test set for evaluation. The test set is disjoint with the generator’s training data for rigorous assessment. To ensure perturbation invisibility, we follow [wang2024transferable] and limit the perturbation budgets to and to . For augmentation, we resize the original images into five scales , and apply Gaussian noise . See Appendix B for more details.
@@ 4.2 Universal Attack Effectiveness
To align with previous studies [zhang2022towards, lu2023set], we first consider the typical V+L task image-text retrieval and calculate the ASR as the proportion of successful adversarial samples within the originally correctly predicted pairs. We present the performance based on R@1 retrieval results in Table 1. Appendix C supplements the results of R@5 and R@10.
@@
White-box attack performance. By observing the white-box ASR in the gray-shaded area, we demonstrate that the proposed algorithm stably achieves excellent ASR on all the evaluated VLP models, validating the outstanding capability of the produced UAP. With only a single pair of perturbations, we reach a noteworthy average white-box ASR of nearly 90% on two large datasets in terms of both TR and IR tasks. Compared with GAP [poursaeed2018generative] and ETU [zhang2024universal], the proposed method consistently enhances the fooling rates in the white-box scenario, confirming the great validity of our suggested multimodal contrastive-learning mechanism. Essentially, the exceptional performance stems from the efficacy of our generated UAP in destroying the alignment between the image and text modalities, thereby misleading the VLP model during inference.